​Assume Evco, Inc. has a current stock price of $49.74 and will pay a $2.20 dividend in one year


Assume Evco, Inc. has a current stock price of $49.74 and will pay a $2.20 dividend in one year, its equity cost of capital is 18%. What price must you expect Evco stock to sell for immediately after the firm pays the dividend in one year to justify its current price? 


We can expect Evco stock to sell for $_______  (Round to the nearest cent.)
Answer #1

Ans: Formula that combines , current price and current selling price.

Po = ( Div1 + P1 ) / (1+rE)

P0 = Current market price per share = $ 49.74

Div1 = $ 2.2

rE = 18%

P1 = New market price = ?

49.74 = (2.2 + P1) / (1+0.18)

49.74 * 1.18 = 2.2+P1

58.6932 = 2.2 + P1

P1 = 56.4932

= $ 56